Corporate Finance

What is the future value of 2,100 in 17 years assuming an interest rate of 8.4 percent compounded semiannually?
Answer #1

C=2100

I=8.4%

n=17

FV = C(1+I)n

=2100(1+0.084/2)17=$4226.34

answered by: Rachale
Answer #2

C=2100


I=8.4%


n=17


FV = C(1+I)n


=2100(1+0.084/2)17/2= the answer


don't forget to divide the exponent, rachale!

answered by: sb